what is the exact value of 6cos(-61pie/6) - 2sin(13pie/4)
Answer by venugopalramana(3286)

what is the exact value of 6cos(-61pie/6) - 2sin(13pie/4)
=6COS(61PI/6)-2SIN{2PI+PI+(PI/4)}
=6COS{2*5PI+(PI/6)}-2SIN{PI+(PI/4)}
=6COS(PI/6)+2SIN(PI/4)
=6*SQRT(3)/2+2*1/SQRT(2)
=3*1.732+SQRT(2)
=5.196+1.414=6.61